# Hongzhi Zang1, Mingjie Wei6,2, Si Xu3, Yongji Wu5, Zhen Guo3, Yuanqing Wang4,3, Hao Lin3, &Liangzhi Shi1, Yuqing Xie1, Zhexuan Xu1, Zhihao Liu7,2, Kang Chen4,2, Wenhao Tang1, &Quanlu Zhang3, Weinan Zhang6, Chao Yu1,2,†,∗, Yu Wang1,†

### v1

@@ 5.3.2 Tips for GRPO
Success rate filtering can improve training stability in some settings of GRPO.
As discussed in Section 4.2.2, we implement a success rate filter for GRPO that discards groups in which all trajectories have identical cumulative rewards. This mechanism can improve the stability of GRPO training. For instance, in the OpenVLA ManiSkill setting, Figure 14(a) shows that training without the filter (“w/o Filter”) exhibits a clear collapse around step 400, whereas enabling the filter (“w/ Filter”) largely alleviates this issue. However, the benefit of the filter is not universal: in the OpenVLA-OFT ManiSkill setting (Figure 14(b)) and the OpenVLA-OFT LIBERO-Goal setting (Figure 14(c)), its effectiveness is much less pronounced.
5.3.3 Hyperparameters Choice
Effect of rollout data size.
@@ 5.3.3 Hyperparameters Choice
(a) PPO, LIBERO-10
(b) GRPO, LIBERO-130
Figure 15: Ablation study on rollout data size. Darker colors correspond to larger rollout datasets. “x Trajs” denotes the number of trajectories included in the rollout data per iteration.
Using LoRA may not directly affect performance but often requires different hyperparameters.
Figure 16 shows the training curves with and without LoRA. The x-axis denotes the number of training epochs, while the y-axis shows the success rate in ManiSkill. The curves are overall similar, suggesting that LoRA itself does not substantially change performance. However, the choice of whether to use LoRA can influence the optimal hyperparameters. For example, in GRPO experiments (Figure 16(b)), using the same learning rate of $1\times 10^{-4}$ leads to normal improvement with LoRA, but the success rate without LoRA collapses to zero. In contrast, when the learning rate is reduced to $1\times 10^{-5}$, the non-LoRA setting also achieves stable improvement. These results indicate that different LoRA configurations may require separate hyperparameter tuning.
@@
Figure 16: Ablation of LoRA in ManiSKill with OpenVLA-OFT model.
5.4 Real-World Deployment
We further evaluate the transferability of simulation-trained models by deploying them to the real world.
Experimental Setup
The hardware setup consists of a Franka Panda robotic arm and an Intel RealSense D435 camera. The relative pose between the camera and the robot is calibrated to ensure consistency with the simulation environment. For real-world evaluation, we compare the “OpenVLA (RLinf-PPO)” model with “OpenVLA (SFT)”. For “OpenVLA (SFT)”, we train OpenVLA to converge on datasets ranging from a few hundred to 64k expert trajectories (1.26M transitions) in simulation. We observe that performance plateaus at roughly 16k trajectories, so we adopt the 16k-trajectory SFT checkpoint as the baseline for comparing RL fine-tuning methods. The evaluation involves six previously unseen objects (pepper, cucumber, banana, kiwi fruit, mangosteen, and sponge), each tasked with being placed into a bowl. For every object, five trials are conducted under different initial configurations of both the object and the bowl.
Table 6: Real-world deployment results. “Pick” indicates that the policy successfully grasped the object, while “Success” indicates that the object was both picked and placed into the bowl.
Object Type OpenVLA (SFT) OpenVLA (RLinf-PPO)
Pick Success Pick Success
Pepper 1/5 0/5 2/5 1/5
Cucumber 1/5 0/5 1/5 1/5
Banana 0/5 0/5 3/5 3/5
Kiwi 0/5 0/5 2/5 1/5
Mangosteen 1/5 0/5 1/5 1/5
Sponge 0/5 0/5 4/5 1/5
Total 3/30 0/30 13/30 8/30
Figure 17: Example of real-world deployment with an RL-trained model performing the task of placing a banana into a bowl.
Results
Table 6 reports the average performance in real-world trials. OpenVLA (RLinf-PPO) substantially outperforms OpenVLA (SFT). Figure 17 illustrates a representative example, where the model successfully places a banana into a bowl. Notably, beyond basic calibration between simulation and the real world, no additional sim-to-real adaptation techniques were employed. Despite the preliminary nature of the experiment, the RL-trained policy demonstrates zero-shot transfer to real-world tasks, revealing the strong generalization capability enabled by simulation-based reinforcement learning. We observe that the SFT policy tends to exhibit excessive motion and overshooting, often resulting in inaccurate grasp poses and task failures. In contrast, the RL-trained policy, though showing minor oscillations, iteratively refines the end-effector pose and achieves a higher grasp success rate.
6 Conclusion
In this work, we introduced RLinf-VLA, a unified and efficient framework for reinforcement learning-based training of Vision-Language-Action models. RLinf-VLA integrates multiple simulators, algorithms, and VLA architectures, while providing flexible execution modes and system-level optimizations that significantly improve training efficiency. Extensive experiments demonstrate that a single unified model achieves state-of-the-art success across a wide range of simulated tasks, and RL-trained policies exhibit stronger zero-shot generalization on a real-world Franka robot compared to supervised fine-tuning. Moreover, our study distills actionable best practices for both PPO and GRPO, guiding future research in RL-based VLA training. By open-sourcing RLinf-VLA with ongoing maintenance, we provide the community with a foundation to accelerate, standardize, and scale research in embodied intelligence.
7 Future Work
Our long-term goal is to advance the capabilities of VLA models through reinforcement learning. Achieving this requires a reliable and extensible infrastructure, which we will continue to actively maintain and expand. Below, we outline several directions for future development.
  • We plan to broaden simulator support. Integration with RoboTwin (Mu et al., 2025) and IsaacLab (Developers, ) is currently in progress and will be released soon, enabling researchers to explore a wider range of embodied environments.
  • Second, we will extend the set of supported models. In particular, integrations with $\pi_{0}$ (Black et al., 2024) and $\pi_{0.5}$ (Intelligence et al., 2025) will be available in the near future, along with algorithmic tuning to ensure strong performance out of the box.
  • Third, we aim to enrich the algorithmic coverage of the framework. Beyond the current on-policy methods, we are working to incorporate classical off-policy algorithms such as Soft Actor-Critic (SAC) (Haarnoja et al., 2018), which are known for their sample efficiency and stability.
  • Finally, while our current focus has been on simulation-based RL, we will also integrate real-world reinforcement learning into the codebase. This step is critical for bridging the gap between simulation and real-world deployment, and will further highlight the practicality of our framework.
In the longer term, we envision this framework as a foundation for establishing standardized benchmarks and fostering community collaboration, thereby accelerating progress in the development and deployment of VLA models.

### v2

@@ X-B Ablation Studies
Success rate filtering
As shown in Figure 14, we implement a success rate filter for GRPO that discards groups in which all trajectories have identical cumulative rewards. This mechanism can improve the stability of GRPO training. For instance, in the OpenVLA ManiSkill setting, training without the filter (“w/o Filter”) exhibits a clear collapse around step 400, whereas enabling the filter (“w/ Filter”) largely alleviates this issue. However, the benefit of the filter is not universal: in the OpenVLA-OFT ManiSkill setting and the OpenVLA-OFT LIBERO-Goal setting, its effectiveness is much less pronounced.
(a) PPO, LIBERO-Long
(b) GRPO, LIBERO 130 tasks
Figure 15: Ablation study on rollout data size. Darker colors correspond to larger rollout datasets.
Effect of rollout data size
@@
Figure 16: Ablation of LoRA in ManiSkill with OpenVLA-OFT.
Using LoRA may not directly affect performance but often requires different hyperparameters
Figure 16 shows the training curves with and without LoRA [13]. The x-axis denotes the number of training epochs, while the y-axis shows the success rate in ManiSkill. The curves are overall similar, suggesting that LoRA itself does not substantially change performance. However, the choice of whether to use LoRA can influence the optimal hyperparameters. For example, in GRPO experiments (Figure 16(b)), using the same learning rate of $1\times 10^{-4}$ leads to normal improvement with LoRA, but the success rate without LoRA collapses to zero. In contrast, when the learning rate is reduced to $1\times 10^{-5}$, the non-LoRA setting also achieves stable improvement. These results indicate that different LoRA configurations may require separate hyperparameter tuning.